i need help with this question
19a-3(a-6)=66
so far i got
19a-3a-18=66

Answer by Isabellaksmith(3)

a = 84/16
19a-3(a-6)=66 .. follow the pyramid formation by completing one thing at a time
19a-3a-18=66..you multiplied everything in the parethesis by 3
16a-18=66..you grouped your a's
16a-18+18=66+18 ..you eliminated your 18 by adding it to each side therfore you are left with :
16a=84 ... divide each side by 16 to leave you with a
a = 84/16 ... or a = 5.25
the revers the pyramid to check ( always check your equations )